OpenStudy (anonymous):

Find the point on the curve x=y^2-8y+ 18 that is closest to the point (-2, 4). Help?

OpenStudy (anonymous):

x - 2 = (y-4)^2 parameterize this parabola

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do u know wht a parabola is?

OpenStudy (anonymous):

yes x^2

OpenStudy (anonymous):

it can be anything of the form which has one variable of deg 1 and another of deg 2 like (x-h) = (y-k)^2

OpenStudy (anonymous):

okay so what does parameterize mean?

OpenStudy (anonymous):

u write x and y in terms of a single variable t

myininaya (myininaya):

d^2=(x-(-2))^2+(y-4)^2 d^2=(y^2-8y+18+2)^2+(y-4)^2 [d^2]'=2(y^2-8y+20)*(2y-8)+2(y-4) set [d^2]'=0 (y^2-8y+20)(2y-8)+(y-4)=0 2(y^2-8y+20)(y-4)+(y-4)=0 (y-4){2(y^2-8y+20)+1}=0 (y-4){2y^2-16y+41}=0 y=4 y={16 + sqrt{16^2-4*2*41}}/4 -> this has imaginary part y={16-sqrt{16^2-4*2*41}}/4-> this has imaginary part so y=4 if y=4, then x=2 so the point (2,4) on the parabola is closest to (-2,4)

OpenStudy (anonymous):

ya those are the steps ty soo much

OpenStudy (anonymous):

im getting old

OpenStudy (anonymous):

he was typing a reply for 10 min^^..knew it was gona be detailed

OpenStudy (anonymous):

she

myininaya (myininaya):

thank you him lol

OpenStudy (anonymous):

you keep on showing ur smarter than me..how old r u?

myininaya (myininaya):

26

OpenStudy (anonymous):

fine then..to accept defeat..im 18

myininaya (myininaya):

lol

myininaya (myininaya):

you still have long way to go you haven't started colllege yet right or first semester?

myininaya (myininaya):

i still have long way to go

OpenStudy (anonymous):

bt weve been taught a lot of this at school..in india..its jujst tht i dont think simple

OpenStudy (anonymous):

what uni/college do you attend?

OpenStudy (anonymous):

oh in india o.o

OpenStudy (anonymous):

going to attend one

OpenStudy (anonymous):

the list is yet to come out

myininaya (myininaya):

I attend UALR its a school in arkansas im going to start my phd if i do well on the gre

OpenStudy (anonymous):

phd on what?

OpenStudy (anonymous):

congratz

myininaya (myininaya):

im going the less math route (i mean it still has math involved, but its narrowed down to computational number theory) im going to do applied science but my researh will be of theory of cryptography

myininaya (myininaya):

i figured out i'm not really a math person i mean to be a math person you have to enjoy all math and i don't

OpenStudy (anonymous):

too advanced for me

myininaya (myininaya):

i still enjoy alot of math though i mean i come on like everyday, but this stuff we do on here is mostly fun its computational its not really math lol

myininaya (myininaya):

do i make sense lol

myininaya (myininaya):

just ignore me lol

OpenStudy (anonymous):

no ur making sense lol

myininaya (myininaya):

so lazy did you understand everything i did up there? i use the formula for distance squared

myininaya (myininaya):

we wanted to minimize the distance so i took derivative and set =0

OpenStudy (anonymous):

i understand the formula but not how you got the derivative in so little steps

OpenStudy (anonymous):

i understand the formula but not how you got the derivative in so little steps

OpenStudy (anonymous):

i understand the formula but not how you got the derivative in so little steps

OpenStudy (anonymous):

i understand the formula but not how you got the derivative in so little steps

myininaya (myininaya):

i used chain rule for both parts

OpenStudy (anonymous):

everything is under a squre root so there should be a 1/2 at the start and the whole thing should be to the -1/2?

myininaya (myininaya):

i see what you are saying don't do the derivative of distance take the derivative of distance squared let f(y)=d^2 and find f'(y)

myininaya (myininaya):

it is easier but lazy you can do derivative of d it will jjust take longer

OpenStudy (anonymous):

hmm interesting , i never learned it this way

myininaya (myininaya):

did you always do it unsquared?

OpenStudy (anonymous):

yes

myininaya (myininaya):

both ways are not wrong just so you know you come to same conclusions both ways minimizing d^2 will give you same result as minimizing d

OpenStudy (anonymous):

how do i take the derivative your way?

myininaya (myininaya):

so you have d^2 above right? d^2=(y^2-8y+20)^2+(y-4)^2 f(y)=(y^2-8y+20)^2+(y-4)^2 f'(y)=2(y^2-8y+20)*(2y-8)+2(y-4) used chain rule

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ty calculus exam tomorrow gota get some sleep. Night , appreciate the help

myininaya (myininaya):

later

OpenStudy (anonymous):

best of luck